OpenStudy (anonymous):

sin^2(x) using de moivre's law

OpenStudy (anonymous):

It has been a while since I last saw complex trig but I think that you can write \( \sin ^2 x \) as a function of \( \cos (2x) \) and then use de Moivre's formula there.

OpenStudy (anonymous):

That is, \[ \sin ^2 x = \frac{1}{2} - \frac{1}{2} cos(2x) \]Then \( \cos (2x) = (2 \cos x)^2 - 1 \)

OpenStudy (anonymous):

how can i go from sin^2x to cos2x